OpenStudy (anonymous):

f(x)= log10^x and g(x) = 3x-1 what is f(x) *g(x)

OpenStudy (anonymous):

@hartnn Can you help me please :)

OpenStudy (campbell_st):

well the way the question is written and using some log laws \[\log 10^x = x\] so f(x) = x then f(x)*g(x) = x(3x -1)

OpenStudy (anonymous):

Ok! you broke it up

OpenStudy (campbell_st):

well given the way the question is written is assumed it was a base 10 log... so I simplified f(x)...

OpenStudy (anonymous):

Oh cause the only thing like that is (3x-1)^x

OpenStudy (anonymous):

or log10x^3x-1

OpenStudy (campbell_st):

ok... I understand the question now its \[f(x) = \log_{10} (x)\] g(x) = 3x -1 find \[f(x) \times g(x)\] which is \[(3x -1) \times \log_{10}(x)\] applying the log law form multiplication you get \[\log_{10}x^{3x -1}\] that's an answer choice...

OpenStudy (anonymous):

Ok im omw to the library...so ill be back in 4 minutes haha thank you so far

OpenStudy (anonymous):

@campbell_st im back

OpenStudy (anonymous):

can you help with 2 more?

OpenStudy (anonymous):

if f(x)= log3(x+1) what is f^-1(2)

OpenStudy (campbell_st):

ok... so to find the inverse write you equation as \[x = \log_{3}(y + 1)\] to find the inverse... make y the subject of the equation above...

OpenStudy (anonymous):

ok so rewrite it?

OpenStudy (campbell_st):

well I've done that whats the opposite of taking a log...?

OpenStudy (anonymous):

anti log

OpenStudy (campbell_st):

well its raising it to the power of the base... \[\log_{a}(b) = x .... then ...a^{\log_{a}(b)} = a^x\] which becomes \[b = a^x\] so this is the log law you need for this problem \[3^{\log_{3}(y + 1)} = 3^x\] so after simplifying you get \[y + 1 = 3^x\] so what's the final equation with y as the subject.

OpenStudy (anonymous):

y=1?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

y= 3^x -1/

OpenStudy (campbell_st):

nearly you forgot a piece \[y = 3^x - 1\]

OpenStudy (campbell_st):

so this is the inverse equation \[f^{-1}(x) = 3^x - 1\]

OpenStudy (campbell_st):

just substitute x = 2 and evaluate for the inverse...

OpenStudy (anonymous):

idk how

OpenStudy (campbell_st):

well just evaluate \[f^{-1}(2) = 3^2 - 1\]

OpenStudy (anonymous):

hm ok how do I get the 2 out of paren

OpenStudy (campbell_st):

you don't what if left hand side is saying is ... whats the value of the equation when x = 2 so you need to find the value of \[3^2 - 1=\]

OpenStudy (anonymous):

9-1 = 8 !